Shift report 20020115 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 15 2002 - 16:04:26 EST

  • Next message: Apache: "Shift report 20020115 16:00-24:00"

    ------------------------------------------------------------
    Date: Tue Jan 15 16:04:25 2002, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#12 pages 63 - 64
    
    Shift Summary:
    ------------------------------------------------------------
    1. QLI in the morning, and We had a hour beam. (14:00-15:00)
    
    2. No achievements. 
       Pedestal & DVM : Run 6512,6513, 6514
       Run 6567 : tried to take a MRS trigger only.
    
    
    3. "Big problem in DAQ..."
       -----------------------
       T2 had a readout problem since yesterday evening.
       TPM2 had a "sync error" from time to time.
          
     - Ramiro tried to reset SC VME from DAQ Control.
     - Ramiro & JH checked the status light on the tpc
       readout boards.
     - Ramiro recycled T2 power supply & VME processor.
     - Ramiro will try to replace "readout board" for T2
       when we are able to get in IR.
    
    
      
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Jan 15 2002 - 16:05:08 EST